jonahneedshelp2121:

I REALLY NEED HELP!!! RSTU is a parallelogram. If m∠TUV = 78° and m∠TVU = 54°, explain how you can find the measure of ∠SRV. Show all steps of your work, and refer to any properties of triangles, parallelograms, or triangle congruency theorems as necessary to justify your response.

jonahneedshelp2121:

jonahneedshelp2121:

wow come on i really need help!!!!!

jonahneedshelp2121:

yaaa typing !!!!

jkmiec26:

So first you would have to find out what m<VTU would be.

jonahneedshelp2121:

ok idk how to find that out

jonahneedshelp2121:

wait

jonahneedshelp2121:

here are there steps in need the explaination part

jonahneedshelp2121:

In SVT, m∠TSV+∠SVT+∠STV=180 [Sum of angles of triangle] 31+126+∠STV=180 ∠STV=23 ∠STV=∠URV [Alternate interior angles on the parallel lines, ST and RU] ∠URV=23

jkmiec26:

So it would be 54+78= m<V & m<U Then 180- m<V & m<U = m<T

jonahneedshelp2121:

hu? can you add it to what i sent then send i? like this form?

jonahneedshelp2121:

In SVT, m∠TSV+∠SVT+∠STV=180 [Sum of angles of triangle] 31+126+∠STV=180 ∠STV=23 ∠STV=∠URV [Alternate interior angles on the parallel lines, ST and RU] ∠URV=23

jonahneedshelp2121:

cuz i am bad at this and don't understand it ):

jonah13:

rlly